# Natural Counterfactuals With Necessary Backtracking

### v2

@@ 5 A Method for Generating Natural Counterfactuals
Notice that the equation could have no solutions, i.e., there is no feasible intervention, when the two constraints impossibly hold at the same time. In practice, the Lagrangian method (Boyd & Vandenberghe, 2004) is used to optimize our objective loss as below:
$\mathcal{L}=\left\|an({\bf A})-an({\bf A})^{*}\right\|_{1}+$ (5)
$w_{\epsilon}\sum_{j}[\max(\epsilon-F({\bf V}_{j}={\bf v}_{j}^{*}|pa_{j}^{*}),0)+\max(\epsilon+F({\bf V}_{j}={\bf v}_{j}^{*}|pa_{j}^{*})-1,0)]$
$*\displaystyle s.t.\quad{\bf A}={\bf a}^{*}$
where the function $\max(\cdot)$ returns the maximum value between two given values. The first term represents the measure of distance between two distinct worlds, while the second term enforces the constraint of $\epsilon$-natural generation. Here, the constant hyperparameter $w_{\epsilon}$ serves to penalize noise values situated in the tails of noise distributions.
6 Case Studies
In this section, we apply our method and evaluate its effectiveness through empirical experiments on four synthetic datasets and two real-world datasets: MorphoMNIST and 3DIdentBOX.
We propose using the difference between generated and actual outcomes as a measure of performance. We expect our natural counterfactuals to significantly reduce error compared to non-backtracking counterfactuals. This advantage can be attributed to the effectiveness of our method in performing necessary backtracking that identifies feasible interventions, keeping counterfactual values within the data distribution, particularly when direct interventions are not feasible. On the other hand, non-backtracking counterfactuals, relying solely on direct interventions, often produce out-of-distribution values, posing challenges for machine learning model generalization.
@@ 6.1 Simulation Experiments
We start with four simulation datasets, which we use designed SCMs to generate. Please refer to the Appendix for more details about these datasets. Let first devolve into Toy 1, which contains three variables $(n_{1},n_{2},n_{3})$. $n_{1}$ is the confounder of $n_{2}$ and $n_{3}$, and $n_{1}$ and $n_{2}$ causes $n_{3}$.
Experimental Settings. Assume data and a causal graph are known, but not the ground-truth SCMs. We employ normalizing flows to capture the causal mechanisms of variables $(n_{1},n_{2},n_{3})$. Given the learned SCMs and a data point from the test set as evidence, we set $n_{1}$ or $n_{2}$ as target variable ${\bf A}$ and randomly sample values from test dataset as counterfactual values of the target variable $n_{1}$ or $n_{2}$. For our natural counterfactuals, we use Eqn. 5 to determine feasible interventions, with $\epsilon=10^{-4}$ and $w_{\epsilon}=10^{4}$. In non-backtracking counterfactuals, $n_{1}$ or $n_{2}$ is directly intervened on. We report the Mean Absolute Error (MAE) between our learned counterfactual outcomes and ground-truth outcomes on $n_{2}$ or/and $n_{3}$ with multiple random seeds. Notice there may be no feasible interventions for some changes, as we have claimed in Sec. 5, and thus we only report outcomes with feasible interventions, which is within the scope of our natural counterfactuals.
Table 1: MAE Results on Toy $1-4$. For simplicity, we use $do$ operator in the table to save room, and when natural counterfactuals are referred to, $do$ means $change$. Our approach exhibits an obvious MAE reduction when applied to $n_{2}$ as expected.
@@
Visualization of Counterfactuals on a Single Sample. We assess the counterfactual outcomes for a sample $(n_{1},n_{2},n_{3})=(-0.59,0.71,-0.37)$, aiming for $change(n_{2}=0.19)$. In Fig. 2 (a), we depict the original data point (yellow), the non-backtracking counterfactual (purple), and the natural counterfactual (green) for $(n_{1},n_{2})$. The ground-truth support for these variables is shown as a blue scatter plot.
(1) Feasible Intervention VS Hard Intervention. Non-backtracking counterfactuals apply a hard intervention on $n_{2}$ ($do(n_{2}=0.19)$), shifting the evidence (yellow) to the post-intervention point (purple), which lies outside the support of $(n_{1},n_{2})$. This implies that direct interventions can result in unnatural values. Conversely, our natural counterfactual (green) remains within the support of $(n_{1},n_{2})$ due to necessary backtracking and feasible intervention on $n_{2}$.
(2) Outcome Error. We calculate the absolute error between $n_{3}$’s model prediction and ground-truth value using either the green or purple point as input for the model $p(n_{3}|n_{1},n_{2})$. The error for the green point is significantly lower at $0.03$, compared to $2.31$ for the purple point. This lower error with the green point is because it stays within the data distribution after a feasible intervention, allowing for better model generalization than the out-of-distribution purple point.
@@ 6.2 MorphoMNIST
(a) Causal Graph
(b) Samples
Figure 3: Causal Graph and samples of Morpho-MNIST.
MorphoMNIST (Castro et al., 2019) involves three variables $(t,i,x)$. As depicted in Fig. 3 (a), $t$ (digit stroke thickness) causes both $i$ (stroke intensity) and $x$ (images), with $i$ being the direct cause of $x$.
Table 2: Ablation Study on $\epsilon$
@@
In our experiments, mirroring those in Section 6.1, we incorporate two key changes. First, we utilize two advanced deep learning models, V-SCM (Pawlowski et al., 2020) and H-SCM (Ribeiro et al., 2023), for doing counterfactuals. Second, due to the absence of ground-truth SCM for assessing outcome error, we adopt the counterfactual effectiveness metric from (Ribeiro et al., 2023), as developed in (Monteiro et al., 2023). This involves training a predictor on the dataset to estimate parent values $(\hat{t},\hat{i})$ from a counterfactual image $x$ generated by learned model $p(x|t,i)$ with the input $(t,i)$, and then computing the absolute error $|t-\hat{t}|$ or $|i-\hat{i}|$ between the input $(t,i)$ and their predicted counterparts $(\hat{t},\hat{i})$ as inferred from counterfactual images.
Ablation Study on Naturalness Threshold $\epsilon$. Table 2 demonstrates that our error decreases with increasing $\epsilon$, regardless of whether V-SCM or H-SCM is used. This trend suggests that a larger $\epsilon$ sets a stricter standard for naturalness in counterfactuals, enhancing the feasibility of interventions and consequently lowering prediction errors. This improvement is likely because deep-learning models are more adept at generalizing to high-frequency data (Engstrom et al., 2019).
(a) Causal Graph
(b) Weak Causal Relationship
(c) Strong Causal Relationship
(a) Results of Non-backtracking Counterfactuals
(b) Results of Natural Counterfactuals
Figure 5: Visualization Results on Stong-3DIdent (View the enlarged figure in Fig. 10 in the Appendix).
Table 3: Results on Weak-3DIdent and Strong-3DIdent (abbreviated as “Weak” “Strong” for simplicity). For clarity, we use ”Non” to denote Nonbacktracking.
@@
In this study, we employ two practical public datasets from 3DIdentBOX (Bizeul et al., 2023), namely Weak-3DIdent and Strong-3DIdent. Both datasets share the same causal graph, as depicted in Fig. 4 (a), which includes an image variable $x$ and its seven parent variables, with three pairs of parent variables: $(h,d)$, $(v,\beta)$, and $(\alpha,\gamma)$, where one is the direct cause of the other in each pair. The primary distinction between Weak-3DIdent and Strong-3DIdent lies in the strength of the causal relationships between each variable pair, with Weak-3DIdent exhibiting weaker connections (Fig. 4 (b)) compared to Strong-3DIdent (Fig. 4 (c)). Our approach mirrors the MorphoMNIST experiments, using H-SCM as the learned SCM with $\epsilon=10^{-3}$.
Influence of Causal Effect Strength. As Table 3 reveals, our method outperforms non-backtracking on both datasets, with a notably larger margin in Strong-3DIdent. This increased superiority is due to a higher incidence of infeasible hard interventions in non-backtracking counterfactuals within the Strong-3DIdent dataset.
See Appendix for More Details. For in-depth information on datasets, experimental results’ standard deviation, model training settings, Feasible Intervention Optimization, differences between our natural counterfactuals and related works, and mechanism distance for natural counterfactuals, refer to the Appendix.
@@ 7 Conclusion
To address the impractical scenarios often generated by non-backtracking counterfactuals, we introduces natural counterfactuals. This concept emphasizes real-world applicability by determining feasible intervention, thereby aligning counterfactuals with the data distribution of the actual world. By integrating a naturalness constraint with necessary backtracking, we have developed an optimization framework designed to determine feasible interventions. This method has been effectively demonstrated through various deep learning case studies, highlighting its capability to yield more dependable and precise counterfactuals. Future work involving natural counterfactuals could focus on exploring their applicability and impact in crucial areas like explainable AI and healthcare.
@@ A.1 Toy Datasets
where there are three endogenous variables $(n_{1},n_{2},n_{3})$ and three noise variables $(u_{1},u_{2},u_{3})$. $n_{1}$ is the confounder of $n_{2}$ and $n_{3}$ and $n_{1}$ and $n_{2}$ causes $n_{3}$.
@@
where there are three endogenous variables $(n_{1},n_{2})$ and three noise variables $(u_{1},u_{2})$. $n_{1}$ causes $n_{2}$.
@@
where there are three endogenous variables $(n_{1},n_{2},n_{3},n_{4})$ and three noise variables $(u_{1},u_{2},u_{3},u_{4})$. $n_{1}$ is the confounder of $n_{1}$ and $n_{4}$. $(n_{2},n_{3},n_{4})$ is a chain, i.e., $n_{2}$ causes $n_{3}$, followed by $n_{4}$.
@@
Table 4: MAE Results on Toy $1-4$. For simplicity, we use $do$ operator in the table to save room, and when natural counterfactuals are referred to, $do$ means $change$.
@@
Visualization Results on Toy 1. In Fig. 7, larger counterfactual images are displayed, which are identical to those shown in Fig. 2, with the only difference being their size.
@@ A.2 MorphoMNIST
Figure 8: Causal Graph and samples of Morpho-MNIST.
The MorphoMNIST comes from (Pawlowski et al., 2020), where there are 60000 images as training set and 10,000 images as test dataset. Fig. 8 (a) shows the causal graph for generating Morpho-MNIST; specifically, stroke thickness $t$ causes the brightness intensity $i$, and both thickness $t$ and intensity $i$ cause the digit $x$. Fig. 8 (b) show some samples from Morpho-MNIST. The ground-truth SCMs to generate the Morpho-MNIST are as follows:
@@
Intersection between Ours and NB (NCO=1, NB=1) (NCO=1, NB=0) (NCO=0, NB=1) (NCO=0, NB=0)
@@
Quantitative Results of $change(i)$ or $do(i)$. We use V-SCM to do counterfactual task of $change(i)$ (where $\epsilon=10^{-3}$) or $do(i)$ with multiple random seeds on test set. In Table 5, the first column shows the MAE of $(t,i)$, indicating our results outperform that of non-backtracking, since our approach consistently determine feasible interventions.
@@
The Effectiveness of Feasible Intervention Optimization (FIO). Next, we focus on the rest four-column results. In both types of counterfactuals, we use the same value $i$ in $do(i)$ and $change(i)$. Hence, after inference, we know which image satisfies $\epsilon$-natural generation in the two types of counterfactuals. In ”NC=1” of the table, NC indicates the set of counterfactuals after feasible intervention optimization. Notice that NC set does not mean the results of natural counterfactuals, since some results do still not satisfy $\epsilon$-natural generation after FIO. “NC=1” mean the set containing data points satisfying $\epsilon$-natural generation and “NC=0” contains data not satisfying $\epsilon$-natural generation after feasible intervention optimization. Similarly, “NB=1” means the set containing data points satisfying naturalness criteria. (NC=1, NB=1) presents the intersection of “NC=1” and “NB=1”. Similar logic is adopted to the other three combinations. The number of counterfactual data points are $10,000$ in two types of counterfactuals.
@@
In (NC=1, NB=1) containing $5865$ data points, our performance is similar to the non-backtracking, showing feasible intervention optimization tends to backtrack as less as possible when hard interventions have satisfied $\epsilon$-natural generation. In (NC=1, NB=0), there are $3159$ data points, which are “unnatural” points in non-backtracking counterfactuals. After natural counterfactual optimation, this huge amount of data points becomes “natural”. Here, our approach significantly reduces errors, achieving a $62.8\%$ reduction in thickness $t$ and $49.8\%$ in intensity $i$, the most substantial improvement among the four sets in Table 5. The number of points in (NC=0, NB=1) is zero, showing the stability of our algorithm since our FIO framework will change the hard, feasible intervention into unfeasible intervention. Two types of counterfactuals perform similarly in the set (NC=0, NB=0), also showing the stability of our approach.
Visualization of Counterfactual Images. Fig. 9 shows counterfactual images (second row), based on the evidence images (first row), with intended changes on $i$. The third row illustrates the differences between evidence and counterfactual images. Focusing on the first counterfactual image from non-backtracking and natural counterfactuals respectively, in non-backtracking, despite $do(i)$ where thickness value $4.2$ should remain unchanged, the counterfactual image shows reduced thickness, consistent with the measured counterfactual thickness of $2.6$. In contrast, natural counterfactuals yield an estimated counterfactual thickness ($t^{*}$ in MS) closely matching original counterfactual thickness ($t^{*}$ in CF), due to backtracking for a feasible intervention on the earlier causal variable $t$, thereby maintaining $(t,i)$ within the data distribution. Observing other images also shows larger errors in non-backtracking counterfactual images.
@@
The 3DIdentBOX datasets, first introduced in Bizeul et al. (2023), come with official code for generating customized versions of these datasets. They consist of images created with Blender, each depicting a teapot with seven attributes, such as position, rotation, and hue, determined by seven ground-truth variables.
In our experiment with the 3DIdentBOX, which comprises six datasets, we focus on the positions-rotations-hue dataset. We expand this into two datasets, Weak-3DIdent and Strong-3DIdent. Each dataset includes seven variables, besides the image variable $x$, with specifics outlined in Table 6. Every image features a teapot, with variables categorized into three groups: positions $(x,y,z)$, rotations $(\gamma,\alpha,\beta)$, and hue $b$, representing seven teapot attributes, as depicted in 11 (a). Fig. 11 (b) illustrates that both datasets share the same causal graph. The distributions of several parent variables of image $x$ in these datasets are detailed in Table 7.
Visualization on Strong-3DIdent. In Fig. 10, larger counterfactual images are displayed, which are identical to those shown in Fig. 5, with the only difference being their size.

### v3

@@ 5 A Practical Method for Generating Natural Counterfactuals
Again, in theory, there may be no solution for $an(A)^{*}$ if the naturalness criterion is demanding. In the extreme case, for example, even when ${\bf a}^{*}={\bf a}$, if $\epsilon$ is set so high that the actual instance does not satisfy the condition of $\epsilon$-natural generation, then no solution exists.
We propose to solve this optimization problem using the following approximate method. Since the estimated functions $\hat{f}_{j}$ are assumed to be reversible, we can reformulate the problem of searching for optimal values of the endogenous variables as one of searching for optimal values of the exogenous noise variables. A feasible approach is to use the Lagrangian method (Boyd & Vandenberghe, 2004) to minimize the following objective loss:
$\mathcal{L}({\bf u}^{*}_{{\mathbf{AN}}({\bf A})})=\sum_{{\bf u}_{j}^{*}\in{\bf u}^{*}_{{\mathbf{AN}}({\bf A})}}|\hat{f}_{j}(pa_{j}^{*},{\bf u}_{j}^{*})-{\bf v}_{j}|+w_{\epsilon}\sum_{{\bf u}_{j}^{*}\in{\bf u}^{*}_{{\mathbf{AN}}({\bf A})}}[\max(\epsilon-F({\bf u}_{j}^{*}),0)+\max(\epsilon+F({\bf u}_{j}^{*})-1,0)]$ (6)
$s.t.\quad{\bf u}^{*}_{\bf A}=\hat{f}^{-1}_{\bf A}(pa^{*}_{\bf A},{\bf a}^{*})$
where the optimization parameters are the counterfactual values of the noise variables corresponding to ${\bf A}$’s ancestors, ${\bf u}^{*}_{{\mathbf{AN}}({\bf A})}$, and the function $\max(\cdot)$ returns the maximum between two values. The first term is the distance measure in the FIO problem, while the second term implements the constraint of $\epsilon$-natural generation. The hyperparameter $w_{\epsilon}$ serves to modulate the penalty imposed on non-natural values. Notice that, in order to ensure the hard constraint ${\bf A}={\bf a}^{*}$, ${\bf A}$’s noise value ${\bf u}^{*}_{\bf A}$ is not optimized explicitly, since the value ${\mathbf{pa}}^{*}_{\bf A}$ is fully determined by ${\bf a}^{*}$ and other noise values. Hence, only noise values other than ${\bf u}_{\bf A}^{*}$ are optimized. Further details are provided in Sec. D.
For simplicity, we have focused on the case of “full evidence” in presenting our framework and method. But it is straightforward to extend the approach to address cases of “partial evidence”, i.e., ${\bf E}\neq{\bf V}$. In such cases, for the generative task, we can simply sample from $p({\bf V}|{\bf E})$, treating it as full evidence, and obtain a natural counterfactual from the resulting sample.
6 Experiments
In this section, we evaluate the effectiveness of our method through empirical experiments on four synthetic datasets and two real-world datasets.
We propose using the deviations between generated and ground-truth outcomes as a measure of performance. We expect our natural counterfactuals to significantly reduce errors compared to non-backtracking counterfactuals. This advantage can be attributed to the effectiveness of our method in performing necessary backtracking that identifies feasible interventions, keeping counterfactual values within the data distribution, so that the learned functions are applicable. On the other hand, non-backtracking counterfactuals often produce out-of-distribution values (Hassanpour & Greiner, 2019; Schölkopf et al., 2021), posing challenges for generalization using the learned functions.
@@ 6.1 Simulation Experiments
We start with four simulation datasets, which we use designed SCMs to generate. Please refer to the Appendix for more details about these datasets. Let’s first look at Toy 1, which contains three variables $(n_{1},n_{2},n_{3})$. $n_{1}$ is the confounder of $n_{2}$ and $n_{3}$, and $n_{1}$ and $n_{2}$ cause $n_{3}$.
Experimental Settings. Again, we assume data and a causal graph are known, but not the ground-truth SCM. We employ normalizing flows to learn a generative model of variables $(n_{1},n_{2},n_{3})$ compatible with the causal order. Given the pretrained causal model and a data point from the test set as evidence, we set $n_{1}$ or $n_{2}$ as ${\bf A}$ and randomly sample values from test dataset as counterfactual values of the target variable $n_{1}$ or $n_{2}$. In our natural counterfactuals, we use Eqn. 6 to determine LBF interventions, with $\epsilon=10^{-4}$ and $w_{\epsilon}=10^{4}$, while in non-backtracking counterfactuals, $n_{1}$ or $n_{2}$ is directly intervened on. We report the Mean Absolute Error (MAE) between our learned counterfactual outcomes and ground-truth outcomes on $n_{2}$ or/and $n_{3}$ with multiple random seeds. Notice there may be no feasible interventions for some changes, as we mentioned in Sec. 5, and thus we only report outcomes with feasible interventions, which are within the scope of our natural counterfactuals.
Table 1: MAE Results on Toy 1 to Toy 4 (Lower MAE is better). To save room, we also write “$do$” for “$change$” for natural counterfactuals.
@@
Visualization of Counterfactuals on a Single Sample. We assess the counterfactual outcomes for a sample $(n_{1},n_{2},n_{3})=(-0.59,0.71,-0.37)$, given the desired alteration $change(n_{2}=0.19)$. For natural counterfactuals, it is necessary to backtrack to $n_{1}$ to realize the change. This step ensures that the pair $(n_{1},n_{2})$ remains within a high-density area of the data distribution. In essence, our intervention targets the composite variable $C=(n_{1},n_{2})$. On the other hand, for non-backtracking counterfactual, the intervention simply modifies $n_{2}$ to 0.19 without adjusting $n_{1}$, making $(n_{1},n_{2})$ out of data support. In Fig. 1 (a), we depict the original data point (yellow), the non-backtracking counterfactual (purple), and the natural counterfactual (green) for $(n_{1},n_{2})$. The ground-truth support for these variables is shown as a blue scatter plot.
(1) Feasible Intervention VS Hard Intervention. Non-backtracking counterfactuals apply a hard intervention on $n_{2}$ ($do(n_{2}=0.19)$), shifting the evidence (yellow) to the post-intervention point (purple), which lies outside the support of $(n_{1},n_{2})$. This shows that direct interventions can result in unnatural values. Conversely, our natural counterfactual (green) remains within the support of $(n_{1},n_{2})$ due to backtracking and the LBF intervention on $(n_{1},n_{2})$.
(2) Outcome Error. We calculate the absolute error between $n_{3}$’s model prediction and ground-truth value using either the green or purple point as input for the model $p(n_{3}|n_{1},n_{2})$. The error for the green point is significantly lower at $0.03$, compared to $2.31$ for the purple point. This lower error with the green point is because it stays within the data distribution after a LBF intervention, allowing for better model generalization than the out-of-distribution purple point.
@@ 6.2 MorphoMNIST
Table 2: Ablation Study on $\epsilon$ (Lower MAE is better)
@@
As depicted in Fig. 4 (a) of Appendix, $t$ (digit stroke thickness) causes both $i$ (stroke intensity) and $x$ (images), with $i$ being the direct cause of $x$ in MorphoMNIST. In this experiment, mirroring those in Section 6.1, we incorporate two key changes. First, we utilize two advanced deep learning models, V-SCM (Pawlowski et al., 2020b) and H-SCM (Ribeiro et al., 2023). Although both models are referred to as “SCM,” they only learn the conditional distributions of endogenous variables and, in theory, do not capture the functional relationships for out-of-distribution inputs. Second, due to the absence of ground-truth SCM for assessing outcome error, we adopt the counterfactual effectiveness metric from (Ribeiro et al., 2023; Monteiro et al., 2023). This involves training a predictor on the dataset to estimate parent values $(\hat{t},\hat{i})$ from a counterfactual image $x$ generated by model $p(x|t,i)$ with the input $(t,i)$, and then computing the absolute error $|t-\hat{t}|$ or $|i-\hat{i}|$.
Ablation Study on Naturalness Threshold $\epsilon$. Table 2 demonstrates that our error decreases with increasing $\epsilon$, regardless of whether V-SCM or H-SCM is used. This trend suggests that a larger $\epsilon$ sets a stricter standard for naturalness in counterfactuals, enhancing the feasibility of interventions and consequently lowering prediction errors. This improvement is due to that deep-learning models are more adept at generalizing to high-frequency data (Engstrom et al., 2019).
Table 3: MAE Results on Weak-3DIdent and Strong-3DIdent (abbreviated as “Weak” “Strong” for simplicity). Lower MAE is better. For clarity, we use “Non” to denote Nonbacktracking.
@@
In this study, we employ two practical public datasets from 3DIdentBOX (Bizeul et al., 2023), namely Weak-3DIdent and Strong-3DIdent, where each image contains a teapot. Both datasets share the same causal graph, as depicted in Fig. 7 (b) of the Appendix, which includes an image variable $x$ and its seven parent variables, with a single variable $b$ and three pairs of variables: $(h,d)$, $(v,\beta)$, and $(\alpha,\gamma)$, where one is the direct cause of the other in each pair. The primary distinction between Weak-3DIdent and Strong-3DIdent lies in the strength of the causal relationships between each variable pair, with Weak-3DIdent exhibiting weaker connections (Fig. 7 (c)) compared to Strong-3DIdent (Fig. 7 (d)). Our approach mirrors the MorphoMNIST experiments, using H-SCM as the pretrained causal model with $\epsilon=10^{-3}$.
Influence of Causal Strength. As Table 3 reveals, our method outperforms non-backtracking on both datasets, with a notably larger margin in Strong-3DIdent. This increased superiority is due to a higher incidence of infeasible hard interventions in non-backtracking counterfactuals within the Strong-3DIdent dataset.
See Appendix for More Details. Please refer to the Appendix for information on datasets, generated samples on MorphoMNIST and 3DIdentBox, standard deviation of results, settings of model training and FIO, differences between our natural counterfactuals and related works, and more.
@@ 7 Conclusion
Given a non-parametric SCM learned from data and a causal graph, non-backtracking counterfactual inference or generation may be highly unreliable because the corresponding non-backtracking intervention can result in a scenario that is far removed from the data based on which the SCM is learned. To address this issue, we have proposed a notion of natural counterfactuals, which incorporates a naturalness constraint and aims to keep the counterfactual supposition within the support of the training data distribution with minimal backtracking. We also developed a practical method for the generation or inference of natural counterfactuals, the effectiveness of which was demonstrated by empirical results.
In case it is not already transparent, we have not shown, nor did we intend to argue, that natural counterfactuals are superior to non-backtracking counterfactuals for all purposes. Non-backtracking counterfactuals, when correctly inferred, have clear advantages in revealing causal relations between events and the effects of specific actions. However, as we showed in this paper, in the context of data-driven counterfactual reasoning, inference or generation of non-backtracking counterfactuals often go astray due to the challenges of out-of-distribution generalization. Although correct information about non-backtracking counterfactuals has considerable action-guiding values, using unreliable information for that purpose is epistemologically and ethically dubious. We intend our framework of natural counterfactuals to mitigate this kind of risk, though it may appear less elegant theoretically.
Our current method is based on the assumption that the learned functions are invertible. One purpose of using the assumption is to ensure the identifiability of natural counterfactuals when one only has access to endogenous variables. If the assumption does not hold, identifiability is not guaranteed. For example, suppose $Y=XU_{1}+U_{2}$ where $Y$ and $X$ are endogenous variables and $U_{1}$ and $U_{2}$ are exogenous noises, then the counterfactual outcome will not be identifiable. However, if we also assume a known distribution of exogenous variables, then our method can be generalized without assuming invertible functions or independent exogenous variables.
Finally, we hasten to reiterate that we do not claim that the particular distance measures and naturalness measures used in this paper are the only choices or among the best. It will be interesting to study and compare alternative implementations in future work.
Acknowledgment
This material is based upon work supported by the RGC of Hong Kong under GRF13602720, NSF Award No. 2229881, AI Institute for Societal Decision Making (AI-SDM), the National Institutes of Health (NIH) under Contract R01HL159805, and grants from Salesforce, Apple Inc., Quris AI, and Florin Court Capital.
@@ A.1 Toy Datasets
where there are three endogenous variables $(n_{1},n_{2},n_{3})$ and three noise variables $(u_{1},u_{2},u_{3})$. $n_{1}$ is the confounder of $n_{2}$ and $n_{3}$. $n_{1}$ and $n_{2}$ cause $n_{3}$.
@@
where there are two endogenous variables $(n_{1},n_{2})$ and two noise variables $(u_{1},u_{2})$. $n_{1}$ causes $n_{2}$.
@@
where there are four endogenous variables $(n_{1},n_{2},n_{3},n_{4})$ and four noise variables $(u_{1},u_{2},u_{3},u_{4})$. $n_{1}$ is the confounder of $n_{2}$ and $n_{4}$. $(n_{2},n_{3},n_{4})$ is a chain, i.e., $n_{2}$ causes $n_{3}$, followed by $n_{4}$.
@@
Table 4: MAE Results on Toy 1 to Toy 4. For simplicity, we use $do$ operator in the table to save room, and when natural counterfactuals are referred to, $do$ means $change$.
@@
Visualization Results on Toy 1. In Fig. 3, larger figures are displayed, which are identical to those shown in Fig. 1, with the only difference being their size.
@@ A.2 MorphoMNIST
Figure 4: Causal Graph and samples of MorphoMNIST.
The MorphoMNIST comes from (Pawlowski et al., 2020b), where there are 60000 images as training set and 10,000 images as test dataset. Fig. 4 (a) shows the causal graph for generating MorphoMNIST; specifically, stroke thickness $t$ causes the brightness intensity $i$, and both thickness $t$ and intensity $i$ cause the digit $x$. Fig. 4 (b) show some samples from MorphoMNIST. The ground-truth SCM is as follows:
@@
Intersection between Ours and NB (NC=1, NB=1) (NC=1, NB=0) (NC=0, NB=1) (NC=0, NB=0)
@@
Quantitative Results of $change(i)$ or $do(i)$. We use V-SCM to do counterfactual task of $change(i)$ (where $\epsilon=10^{-3}$) or $do(i)$ with multiple random seeds on test set. In Table 5, the first column shows the MAE of $(t,i)$, indicating our results outperform that of non-backtracking, since our approach consistently determine LBF interventions.
@@
The Effectiveness of FIO. Next, we focus on the rest four-column results. In both types of counterfactuals, we use the same value $i$ in $do(i)$ and $change(i)$. We can calculate which image satisfies $\epsilon$-natural generation. In the table, “NC” indicates the set of counterfactuals after FIO. Notice that “NC” set does not mean the results of natural counterfactuals, since some results do still not satisfy $\epsilon$-natural generation after FIO. “NC=1” mean the set containing data points satisfying $\epsilon$-natural generation and “NC=0” contains data not satisfying $\epsilon$-natural generation after FIO. Similarly, “NB=1” means the set containing data points satisfying naturalness criteria in nonbacktracking counterfactuals. (NC=1, NB=1) presents the intersection of “NC=1” and “NB=1”. Similar logic is adopted to the other three combinations. The number of counterfactual data points are $10,000$ in two types of counterfactuals.
@@
In (NC=1, NB=1) containing $5865$ data points, our performance is similar to the non-backtracking, since FIO does not do backtracking when hard interventions have satisfied $\epsilon$-natural generation. In (NC=1, NB=0), there are $3159$ data points, which are “unnatural” points in non-backtracking counterfactuals. After FIO, this huge amount of data points becomes “natural”. Here, our approach significantly reduces errors, achieving a $62.8\%$ reduction in thickness $t$ and $49.8\%$ in intensity $i$, the most substantial improvement among the four sets in Table 5. The number of points in (NC=0, NB=1) is zero, showing the stability of our algorithm since our FIO framework will change the hard, feasible intervention into unfeasible intervention. Two types of counterfactuals perform similarly in the set (NC=0, NB=0), also showing the stability of our approach.
Table 6: Unfeasible solutions per 10,000 instances on MorphoMNIST
$\epsilon$ Unfeasible Solutions
1e-4 794
1e-3 975
1e-2 1166
Ablation Study on $\epsilon$. In Table 6 provided, we report the frequency of unfeasible solutions per 10,000 instances on MorphoMNIST, following optimization within the V-SCM framework. The data reveals a consistent trend: as the value of $\epsilon$ increases, the frequency of unfeasible solutions also rises. This pattern occurs because a higher $\epsilon$ corresponds to a stricter standard of naturalness, making it more challenging to achieve feasible outcomes.
@@
Visualization of Counterfactual Images. Fig. 5 shows counterfactual images (second row), based on the evidence images (first row), with intended changes on $i$. The third row illustrates the differences between evidence and counterfactual images. Focusing on the first counterfactual image from non-backtracking and natural counterfactuals respectively, in non-backtracking, despite $do(i)$ where thickness value $4.2$ should remain unchanged, the counterfactual image shows reduced thickness, consistent with the measured counterfactual thickness of $2.6$. In contrast, natural counterfactuals yield an estimated counterfactual thickness ($t^{*}$ in MS) closely matching original counterfactual thickness ($t^{*}$ in CF), due to backtracking for a LBF intervention on the earlier causal variable $t$, thereby maintaining $(t,i)$ within the data distribution. Observing other images also shows larger errors in non-backtracking counterfactual images.
@@
The 3DIdentBOX datasets, first introduced in (Bizeul et al., 2023), come with official code for generating customized versions of these datasets. They consist of images created with Blender, each depicting a teapot with seven attributes, such as position, rotation, and hue, determined by seven ground-truth variables.
In our experiment with the 3DIdentBOX, which comprises six datasets, we focus on the positions-rotations-hue dataset. We expand this into two datasets, Weak-3DIdent and Strong-3DIdent. Each dataset includes seven variables, besides the image variable $x$, with specifics outlined in Table 7. Every image features a teapot, with variables categorized into three groups: positions $(x,y,z)$, rotations $(\gamma,\alpha,\beta)$, and hue $b$, representing seven teapot attributes, as depicted in 7 (a). Fig. 7 (b) illustrates that both datasets share the same causal graph. It is important to note a distinction between Weak-3DIdent and Strong-3DIdent. In Weak-3DIdent, there exists a weak causal relationship between the variables of each pair, as shown in Fig. 7 (c), whereas in Strong-3DIdent, the causal relationship is stronger, as depicted in Fig. 7 (d). The distributions of several parent variables of image $x$ in these datasets are detailed in Table 8.